OpenStudy (anonymous):

Find the slope of the line passing through the points (-1,-9) and (2,11).

OpenStudy (anonymous):

m(slope) = (y2-y1)/(x2-x1) suppose (-1,-9) is the (x1,y1) and (2,11) is the (x2,y2). then just plug the values of those x and y m = (11-(-9)) / (2-(-1)) = 20/3

OpenStudy (amistre64):

2,11 1,9 ---- 3,20 = 20/3

OpenStudy (anonymous):

Ohh, ok thanks.

OpenStudy (anonymous):

So its 6.66?

OpenStudy (anonymous):

yes, but usually 6.66666... is rounded to 6.67

OpenStudy (amistre64):

the exact value = 20/3 .... but if you want to do a decimal approximation thats good